A solid shaft is to transmit 800kw at 300rpm. if the shaft is not to twist more than 1 degree on a length of 20 diameters and the shear stress is not to exceed 50mn/m2 with a modulus of rigidity of 90gn/m2, design the shaft by determining a suitable diameter


1
Expert's answer
2021-05-13T07:19:43-0400

Given:

"\text{Power}(P)=800kW"

RPM(N)=300rpm

Maximum angle of twist "(\theta)" ="1^0=\frac{\pi}{180}=0.01745\: \text{radians}"

Length (L)=20"\times \text{diameter}(d)"

maximum shear stress("\tau_s" )=50N/mm2

Modulus of rigidity(G)=90 G N/mm2

To find : Suitable diameter of shaft

Formula : Torsional equation

"\frac{\text{Torque}(T)}{\text{Polar M I}(I_p)}=\frac{\tau_s}{\text{Radius}(R)}=\frac{G\times \theta}{L}"

Power(P) ="\frac{2\pi \times N\times T}{60}"


Calculation:

Power(P) ="\frac{2\pi \times N\times T}{60}"


800x103="\frac{2\pi \times 300\times T}{60}"

T=25464.79 N-m

For solid shaft considering strength of shaft,

"T=\frac{\pi}{16}\times \tau_s\times d^3"

After substituting respective values , we get

Diameter(d)=137.40 mm=138 mm (approximately)

Considering angle of twist of shaft,

"T=\frac{\pi}{16}\times \frac{G\theta}{L}\times d^3"

"T=\frac{\pi}{16}\times \frac{90\times1000\times 0.01745}{20d}\times d^3"

"25464.79\times 10^3=\frac{\pi}{32}\times \frac{90\times1000\times 0.01745}{20d}\times d^4"

d=148.93 mm

Considering larger diameter as suitable diameter of shaft(d)=149 mm